# On‑Device LLM for Context‑Aware Wi‑Fi Roaming

### v1

@@ Abstract
Wireless roaming is a critical yet challenging task for maintaining seamless connectivity in dynamic mobile environments. Conventional threshold‑based or heuristic schemes often fail, leading to either sticky or excessive handovers. We introduce the first cross‑layer use of an on‑device large language model (LLM): high‑level reasoning in the application layer that issues real‑time actions executed in the PHY/MAC stack. The LLM addresses two tasks: (i) context‑aware AP selection, where structured prompts fuse environmental cues (e.g., location, time) to choose the best BSSID; and (ii) dynamic threshold adjustment, where the model adaptively decides when to roam. To satisfy the tight latency and resource budgets of edge hardware, we apply a suite of optimizations—chain‑of‑thought prompting, parameter‑efficient fine‑tuning, and quantization. Experiments on indoor and outdoor datasets show that our approach surpasses legacy heuristics and DRL baselines, achieving a strong balance between roaming stability and signal quality. These findings underscore the promise of application‑layer LLM reasoning for lower‑layer wireless control in future edge systems.
@@ 1 Introduction
Wi-Fi roaming is a critical operation for maintaining seamless wireless connectivity as users move through physical environments. Traditionally, roaming logic has assumed relatively stable topologies—stationary access points (APs) and predictable client mobility. However, emerging scenarios are rapidly redefining this assumption. Automotive Wi‑Fi, non‑terrestrial networks (NTN), and device‑to‑device (D2D) “mobile‑AP” scenarios all feature mobility on both the client and AP sides, creating highly dynamic topologies that demand instantaneous, context‑aware decisions at the wireless edge.
@@
Figure 1: Rule‑based handover (legacy) versus on‑device LLM context‑aware handover.
@@ 2. T2)
  • Cross‑layer control via LLM. We demonstrate the first on‑device LLM that reasons in the application layer while issuing real‑time actions in the PHY/MAC stack.
@@
  • Comprehensive evaluation. Indoor and outdoor experiments show our approach outperforms heuristic and DRL baselines, striking a strong trade‑off between roaming stability and link quality
  • Design insights. We distill practical guidelines for deploying LLMs as edge‑level wireless controllers (§5).
@@ 2.1 Wi-Fi Roaming and Mobility Management
In Wi-Fi networks, roaming is the process by which a client device (e.g., a smartphone or laptop) switches between nearby APs to maintain strong connectivity as the user moves. These APs typically belong to the same network and cover overlapping areas (e.g., ESS; extended service set). For example, as a user walks through a building, the device must decide when to disconnect from one AP and connect to another to avoid degraded signal quality or dropped connections.
Roaming is a Layer 2 operation governed by the IEEE 802.11 standard. It plays a critical role in ensuring seamless connectivity in dynamic environments such as offices, campuses, or transit systems.
Most client devices monitor the received signal strength indicator (RSSI) from their current AP and initiate a handover when the signal drops below a predefined threshold, commonly called scanRSSI. For instance, some devices trigger roaming when RSSI falls below $-70$ dBm. Once this threshold is crossed, the device actively scans for candidate APs and evaluates them based on signal strength and additional factors such as channel congestion or PHY capabilities. A new AP is selected if it provides significantly better link quality—typically determined by a relative RSSI improvement (e.g., $+8$ dB while transmitting, or $+12$ dB while idle).
@@
Figure 2: The two failure modes of heuristic Wi‑Fi roaming.
@@
While such heuristics offer simplicity, fixed RSSI thresholds and rule‑based logic often fall short in practice. As illustrated in Fig. 2, poorly tuned thresholds and static rules can result in:
  • Sticky handovers: Staying connected to a weakening AP, leading to degraded throughput and increased latency.
  • Excessive handovers: Switching too frequently between APs, introducing instability and protocol overhead.
To evaluate roaming behavior, we focus on two performance metrics:
AvgRSSI $=\frac{1}{T}\sum_{t=1}^{T}\text{RSSI}_{t},$ (1)
# HO $=\sum_{t=2}^{T}\mathbb{I}[\text{BSSID}_{t}\neq\text{BSSID}_{t-1}],$ (2)
where $T$ is the number of time steps, $\text{RSSI}_{t}$ is the received signal strength at time $t$, and $\text{BSSID}_{t}$ is the AP the device is associated with. AvgRSSI reflects connection quality, while # HO captures roaming stability. These two objectives often conflict: reducing handovers may hurt signal quality, and maximizing RSSI may cause unnecessary roaming. A well-designed roaming policy must intelligently balance these competing goals based on context.
@@ 3.2 Approach: LLM-based Decision with Contextual Information
While CoT improves zero-/few-shot reasoning, we still need domain adaptation via low-rank apaptation (LoRA) (Hu et al., 2022). Then, to align the model’s decisions with user preferences, we apply direct preference optimization (DPO) (Rafailov et al., 2023).
Fine-Tuning. We fine-tune the pre-trained LLM on real Wi-Fi roaming log data. We adopt LoRA for parameter-efficient fine-tuning (PEFT). Using LoRA, we inject small adaptation parameters and even fine-tune on a quantized model to save memory. This adaptation trains the LLM to understand WLAN-specific cues (e.g., what RSSI patterns precede a disconnection) and to output the best AP choice accordingly. The lightweight nature of LoRA fine-tuning allows us to iterate and improve the model without needing enormous compute resources.
@@ 3.3 Evaluation
We evaluate our Task (1): Context-Aware AP Choice using various Wi-Fi roaming scenarios. Two primary metrics are used: (1) AvgRSSI – the average RSSI of the STA over the test sequence, and (2) # HO – the total number of handovers triggered during the test. While other metrics such as throughput or latency may be affected by handover behavior, we focus on the average RSSI and number of handovers, which capture the key trade-off between connection quality and roaming stability.
@@
Detailed experimental setups and parameter settings are provided in Table 7 in the Appendix.
@@ 4.1 Problem Statement: Dynamic Threshold Selection
A major challenge in WLAN roaming is determining when to initiate handover. This is typically controlled by a fixed RSSI threshold (scanRSSI), below which STA begins scanning for alternative APs. However, as previously discussed, a single static value can lead to two undesirable outcomes: setting it too high causes overly frequent scans and roaming (wasting energy and risking ping-pong effects), whereas setting it too low risks lingering on a deteriorating link until performance degrades or the connection drops. Building on the context-aware approach in Section 3, we now aim to dynamically select the threshold in real time, leveraging an on-device LLM.
Yet deploying a large LLM directly on mobile or edge platforms is challenging due to hardware constraints such as limited memory, battery capacity, and compute power. Accordingly, one need to find a way to reduce the LLM’s computation overhead while preserving strong performance in threshold‐selection decisions.
@@ 4.2 Approach: Adaptive Threshold via On-Device LLM
Model Selection and Quantization. Deploying large models on resource-constrained edge devices necessitates careful consideration of model size and computational demands. To ensure practical lower-layer operation, we select a model size that optimally balances inference speed with decision accuracy. We further reduce computational overhead by employing quantization techniques (e.g., Q2_K), significantly compressing model weights without a substantial loss in performance.
@@ 4.3 Experiments
Unless otherwise noted, we evaluate this approach under the same general setup, baselines, and metrics used for Task (1) in Sec. 3.3.
Effect of Quantization. Table 4 illustrates how various quantization schemes (e.g., Q2_K, Q3_K_M, Q4_K_M) reduce the model’s size (from 8.5 GB to as low as 3.2 GB) while keeping roaming performance nearly the same ($\approx$138–145 handovers, $\approx-58$ dBm RSSI). Consequently, we adopt Q2_K in subsequent experiments for its substantial resource savings with minimal accuracy loss.
@@
Table 4: Impact of quantization on roaming performance and model size (Llama3.1-8B) (Grattafiori et al., 2024).
@@
In contrast, the proposed context-aware on-device LLM dynamically adjust its threshold to local context information, enabling it to maintain balanced RSSI performance and moderate handover frequency across diverse scenarios. While it does not always achieve the absolute best performance on every single metric, the LLM notably delivers robust, generalized performance without manual tuning. This underscores the practical value and effectiveness of cross-layer, adaptive threshold optimization using on-device LLMs for real-world WLAN roaming scenarios.
@@ 5 Discussion
In conclusion, this section discuss key insights and practical considerations for deploying LLMs as edge-level decision-making agents for wireless lower-layer operations through the following key questions:
@@
Q3: Is an on-device LLM practical for lower-layer operations?. As discussed in Sec. 3.5, current on-device LLM implementations exhibit inference latencies on the order of seconds, presenting challenges for strict near-real-time (RT) operations. However, for tasks like threshold adjustment—which require decisions less frequently and tolerate slightly higher latency—on-device LLMs demonstrate significant practical feasibility. Importantly, while the LLM resides at the application layer, its decisions directly manipulate parameters at the MAC layer, exemplifying a compelling case of application-layer AI enabling adaptive lower-layer wireless control. Demonstrations and detailed implementation information supporting this practicality are available at github.com/abman23/XXXXX.
@@
Q4: What are the main challenges and future directions?. Deploying on-device LLMs presents several critical challenges, particularly managing computational overhead and optimizing models for strict hardware constraints inherent to edge devices. Although current on-device LLMs show promising capabilities, their universal practicality for all lower-layer wireless operations—especially those demanding strict real-time responsiveness—remains limited. Future research must focus on refining open-source LLM architectures and techniques, emphasizing efficiency improvements in latency, memory usage, and power consumption. Additionally, investigating the broader applicability of on-device LLMs to other essential wireless control tasks beyond WLAN roaming will be crucial in fully harnessing their potential as versatile and practical AI-based decision makers in wireless communication systems.

### v2

@@ Abstract
Roaming in Wireless LAN (Wi-Fi) is a critical yet challenging task for maintaining seamless connectivity in dynamic mobile environments. Conventional threshold‑based or heuristic schemes often fail, leading to either sticky or excessive handovers. We introduce the first cross‑layer use of an on‑device large language model (LLM): high‑level reasoning in the application layer that issues real‑time actions executed in the PHY/MAC stack. The LLM addresses two tasks: (i) context‑aware AP selection, where structured prompts fuse environmental cues (e.g., location, time) to choose the best BSSID; and (ii) dynamic threshold adjustment, where the model adaptively decides when to roam. To satisfy the tight latency and resource budgets of edge hardware, we apply a suite of optimizations—chain‑of‑thought prompting, parameter‑efficient fine‑tuning, and quantization. Experiments on indoor and outdoor datasets show that our approach surpasses legacy heuristics and DRL baselines, achieving a strong balance between roaming stability and signal quality. These findings underscore the promise of application‑layer LLM reasoning for lower‑layer wireless control in future edge systems.
@@ 1 Introduction
Wi-Fi roaming is a critical operation for maintaining seamless wireless connectivity as users move through physical environments. Traditionally, roaming logic has assumed relatively stable topologies—stationary access points (APs) and predictable client mobility. However, emerging scenarios are rapidly redefining this assumption. Automotive Wi‑Fi and device‑to‑device (D2D) “mobile‑AP” scenarios all feature mobility on both the client and AP sides, creating highly dynamic topologies that demand instantaneous, context‑aware decisions at the wireless edge.
@@
Figure 1: Cross-layer control via on-device LLM: Rule‑based handover (legacy) vs. on‑device LLM context‑aware handover.
@@ 2. T2)
  • Cross‑layer Wireless Control via On-Device LLM. We demonstrate the first on‑device LLM that reasons in the application layer while issuing real‑time actions in the PHY/MAC stack.
@@
  • Comprehensive evaluation. Indoor and outdoor experiments show our approach outperforms legacy and DRL baselines, striking a strong trade‑off between roaming stability and link quality
  • Practical insights & Real-world demo. We distill practical guidelines for deploying LLMs as edge‑level wireless controllers (§5) and validate our approach through practical demonstrations (§A) [^1] A full demonstration video and inference code are available at github.com/abman23/on-device-llm-wifi-roaming..
@@ 2.1 Wi-Fi Roaming and Mobility Management
In Wi-Fi networks, roaming refers to the process by which a client device (e.g., a smartphone or laptop) switches its connection from one AP to another as the user moves. These APs typically belong to the same extended service set (ESS), providing overlapping coverage areas. For instance, as a user walks through a building, the device must determine when it should disconnect from the current AP and connect to another AP to maintain a strong wireless connection and avoid dropped connections or degraded signal quality.
Roaming is a Layer 2 operation governed by the IEEE 802.11 standard, playing a crucial role in ensuring seamless connectivity across dynamic environments such as office buildings, university campuses, and public transit stations.
Most client devices continuously monitor the received signal strength indicator (RSSI) of their currently associated AP. When the RSSI falls below a predefined threshold, known as scanRSSI (typically set around $-70$ dBm), the device initiates a roaming procedure. After crossing this threshold, the device actively scans for candidate APs, evaluating them based on factors such as RSSI, channel congestion, and PHY-layer capabilities. It then selects a new AP if this AP offers significantly better link quality, commonly defined by a relative RSSI improvement threshold (e.g., at least $+8$ dB during active data transmission or $+12$ dB during idle periods).
@@
Figure 2: Common failure modes of rule-based Wi-Fi roaming.
@@
Despite their simplicity, fixed RSSI thresholds and rule-based roaming logic frequently lead to suboptimal outcomes. As shown in Fig. 2, poorly tuned thresholds and static decision rules often produce two problematic scenarios:
  • Sticky handovers: The device remains connected to a weakening AP, causing deteriorating throughput, increased latency, and poor user experience.
  • Excessive handovers: The device frequently initiates scanning and switches between APs, causing disruptions in the connection, increased signaling overhead, lowered throughput, and degraded user experience.
Effectively managing roaming thus requires balancing two conflicting objectives: maintaining high signal quality while minimizing handover frequency. To quantitatively evaluate roaming strategies, we consider two key metrics:
  • AvgRSSI: The average RSSI experienced by the client (Station, STA) throughout the test period, reflecting the overall connection quality (higher is better).
  • #HO: The total number of handovers triggered during the evaluation period, indicating roaming stability or the overhead associated with frequent AP switching (lower is better).
A well-designed roaming policy must intelligently and dynamically balance these two metrics based on real-time environmental context and user requirements.
@@ 3.2 Approach: LLM-based Decision with Contextual Information
While CoT improves zero-/few-shot reasoning, we still need domain adaptation via supervised fine-tuning. Then, to align the model’s decisions with user preferences, we apply direct preference optimization (DPO) (Rafailov et al., 2023).
Fine-Tuning. We fine-tune the pre-trained LLM on real Wi-Fi roaming log data. We adopt LoRA (Hu et al., 2022) for parameter-efficient fine-tuning (PEFT). Using LoRA, we inject small adaptation parameters and even fine-tune on a quantized model to save memory. This adaptation trains the LLM to understand Wi-Fi-specific cues (e.g., what RSSI patterns precede a disconnection) and to output the best AP choice accordingly. The lightweight nature of LoRA fine-tuning allows us to iterate and improve the model without needing enormous compute resources.
@@ 3.3 Evaluation
We evaluate the Task (1): Context-Aware AP Choice by comparing our LLM-based approach against several baseline roaming strategies in a variety of Wi-Fi roaming scenarios. To evaluate roaming behavior, we use two key metrics:
AvgRSSI $=\frac{1}{T}\sum_{t=1}^{T}\text{RSSI}_{t},$ (1)
# HO $=\sum_{t=2}^{T}\mathbb{I}[\text{BSSID}_{t}\neq\text{BSSID}_{t-1}],$ (2)
Here, $T$ denotes the number of time steps, $\text{RSSI}_{t}$ is the received signal strength at time $t$, and $\text{BSSID}_{t}$ is the AP the device is associated with. AvgRSSI reflects connection quality, while # HO captures roaming stability. These two objectives often conflict: reducing handovers may hurt signal quality, and maximizing RSSI may cause unnecessary roaming. A well-designed roaming policy must intelligently balance these competing goals based on context.
@@
Detailed experimental setups and parameter settings are provided in Table 7 in Appendix C.
@@ 4.1 Problem Statement: Dynamic Threshold Selection
A major challenge in Wi-Fi roaming is determining when to initiate handover. This is typically controlled by a fixed RSSI threshold (scanRSSI), below which STA begins scanning for alternative APs. However, as previously discussed, a single static value can lead to two undesirable outcomes: setting it too high causes overly frequent scans and roaming (wasting energy and risking ping-pong effects), whereas setting it too low risks lingering on a deteriorating link until performance degrades or the connection drops. Building on the context-aware approach in Section 3, we now aim to dynamically select the threshold in real time, leveraging an on-device LLM.
Yet deploying a large LLM directly on mobile or edge platforms is challenging due to hardware constraints such as limited memory, battery capacity, and compute power. Accordingly, one needs to find a way to reduce the LLM’s computation overhead while preserving strong performance in threshold‐selection decisions.
@@ 4.2 Approach: Adaptive Threshold via On-Device LLM
Model Selection and Quantization. Deploying large models on resource-constrained edge devices requires careful consideration of model size and computational resources. To ensure practical lower-layer operation, we select a model size that optimally balances inference speed with decision accuracy. We further reduce computational overhead by employing the Q2_K quantization scheme. Specifically, Q2_K compresses weights into groups of $16$, encoding each weight with $2$-bit precision and using one shared $4$-bit scale and offset per block. This compact encoding needs only $2.56$ bits per parameter—about 8×8$\times$8 × fewer bits than FP16—while staying fully compatible with the GGUF model format supported by llama.cpp.
@@ 4.3 Experiments
Unless otherwise noted, we evaluate this approach under the same general setup, baselines, and metrics used for Task (1) in Sec. 3.3. Detailed experimental setups and parameter settings are provided in Table 7 in Appendix C.
Effect of Quantization. Q4_K_MTable 4 compares the impact of various quantization schemes (e.g., Q2_K, Q3_K_M, Q4_K_M) on model size and roaming performance. Quantization significantly reduces the model’s footprint from 8.5,GB (Q8_0) down to as low as 3.2,GB (Q2_K), while maintaining comparable performance metrics (approximately 138145absent138145$\approx$ 138–145≈ 138 – 145 handovers and around $\approx-58$,dBm RSSI). Given these results, we select Q2_K quantization for subsequent experiments, as it provides substantial resource savings with negligible loss in decision accuracy.
@@
Table 4: Comparison of quantization schemes for the Llama3.1-8B model in terms of roaming decision and memory footprint (Grattafiori et al., 2024).
The adoption of Q2_K reduces the model size by approximately 2.7×, from 8.5 GB with Q8_0 quantization to 3.2 GB, enabling efficient operation within the 16,GB unified memory capacity of the Apple M-series MacBook used for our on-device experiments.
@@
In contrast, the proposed context-aware on-device LLM dynamically adjust its threshold to local context information, enabling it to maintain balanced RSSI performance and moderate handover frequency across diverse scenarios. While it does not always achieve the absolute best performance on every single metric, the LLM notably delivers robust, generalized performance without manual tuning. This underscores the practical value and effectiveness of cross-layer, adaptive threshold optimization using on-device LLMs for real-world Wi-Fi roaming scenarios.
@@ 5 Discussion
In conclusion, this section discusses key insights and practical considerations for deploying LLMs as edge-level decision-making agents for wireless lower-layer operations through the following key questions:
@@
Q3: Is an on-device LLM practical for lower-layer operations?. As discussed in Sec. 3.5, current on-device LLM implementations exhibit inference latencies on the order of seconds, posing challenges for strict near-real-time (RT) operations. Nevertheless, recent advances in dedicated AI accelerator hardware are expected to significantly reduce inference times, greatly enhancing the practicality of on-device LLMs even for near-RT Wi-Fi tasks such as roaming. Moreover, for scenarios like threshold adjustment—where decisions are required less frequently and slightly higher latencies can be tolerated—the existing on-device LLM deployments already demonstrate substantial feasibility. Importantly, although inference occurs at the application layer, the decisions made directly adjust MAC-layer parameters, clearly illustrating the viability of application-layer AI enabling adaptive lower-layer wireless control. Demonstrations and details are available at github.com/abman23/on-device-llm-wifi-roaming.
@@
Q4: What are the main challenges and future directions?. Deploying on-device LLMs presents several critical challenges, particularly managing computational overhead and optimizing models for strict hardware constraints inherent to edge devices. Although current on-device LLMs show promising capabilities, their universal practicality for all lower-layer wireless operations—especially those demanding strict real-time responsiveness—remains limited. Future research must focus on refining open-source LLM architectures and techniques, emphasizing efficiency improvements in latency, memory usage, and power consumption. Additionally, investigating the broader applicability of on-device LLMs to other essential wireless control tasks beyond Wi-Fi roaming will be crucial in fully harnessing their potential as versatile and practical AI-based decision makers in wireless communication systems.
Appendix A Real-World Demonstration
To validate on-device deployment, we ran our LLM-based roaming agent on a MacBook Pro (Apple M-series, 16 GB RAM, macOS). Figure 8 captures key moments from a live screen-recorded session: the sensing phase (RSSI and context collection) followed by the LLM’s reasoning and threshold update, shown for both indoor and outdoor walks. Comprehensive performance metrics confirm the feasibility of real-time operation on consumer-grade hardware. A full demonstration video is available at github.com/abman23/on-device-llm-wifi-roaming.
(a) Indoor — sensing phase
(b) Indoor — reasoning & decision
(c) Outdoor — sensing phase
(d) Outdoor — reasoning & decision
Figure 8: Real-world demonstration of our on-device LLM for Task 2 (dynamic threshold adjustment). In each setting the laptop (Macbook Pro) first senses RSSI and context information (left), then the LLM reasons and outputs an adaptive roaming threshold (right), validating practical operation indoors and outdoors.
Appendix B Qualitative Assessment of On-Device LLM Reasoning
Figure 9: Reasoning output of the on-device LLM during an indoor roaming scenario. The model assesses AP candidates by eliminating those with declining RSSI values and prioritizes APs with stable signal strength, ultimately selecting BSSID 34:3A:20:79:C8:B2 for handover.
In the real-world indoor demonstration shown in Fig. 9, the on-device LLM demonstrates reasoning aligned with practical wireless network management strategies. The inference process reflects logical heuristics such as filtering out APs exhibiting decreasing signal strength and prioritizing stable or improving RSSI, consistent with best practices for effective roaming decisions. This reasoning is derived purely from local contextual (e.g., timestamp, location) and Wi-Fi measurement (e.g., RSSI-BSSID pairs) information. The correctness of the LLM’s reasoning is validated through the subsequent successful handover, evidenced by system logs. This qualitative evaluation complements the quantitative results presented earlier (§4), reinforcing the capability of on-device LLMs to execute valid, interpretable reasoning suitable for lower-layer wireless control at edge-device.
Appendix C Setup
Table 7 details the hardware, dataset statistics, and model-tuning parameters used in Task 1 (AP selection) and Task 2 (threshold adjustment). All real-world experiments ran on a consumer-grade MacBook Pro with an Apple M-series processor (16 GB RAM, macOS), demonstrating that our on-device LLM pipeline is feasible on typical edge hardware.
HW / SW
Device MacBook Pro (M3, 16GB)
OS macOS
Wi-Fi Specification IEEE 802.11ax (Wi-Fi 6E)
Wi-Fi Library CoreWLAN
Dataset (Wi-Fi Log)
Scenarios Indoor, Outdoor
Time Stamps per Scenario 1944, 5808, 1944
Sampling Interval 1 sec
Training : Test Split 80% : 20%
Task Configuration
scanRSSI Threshold -70 dBm
Contextual Inputs Location + Time
Logs per Handover Decision 10
Threshold Adjustment Interval (Task 2) 30 sec
On-Device LLM Configuration
Base Model Llama-3.1-8B-Instruct
Quantization Scheme Q2_K
Fine-Tuning Configuration
Learning Rate $2\times 10^{-4}$
Batch Size / Grad. Accumulation 4 / 4
Number of Epochs 1
Weight Decay 0.01
LoRA Rank / Alpha 128 / 128
Optimizer AdamW
Table 7: Experimental setup and parameter settings for Tasks 1 & 2.
Appendix D Data Collection
We collected Wi-Fi roaming datasets from diverse indoor and outdoor environments, as illustrated in Fig. 10. Each of the ten sessions recorded approximately $\sim\!1{,}000$ consecutive data points, captured at one-second intervals. Each data point captures detailed information, including:
  • AP scan: BSSID, RSSI, etc.
  • Device context: timestamp, geographical coordinates (latitude and longitude), battery status, etc.
The resulting dataset underpins both post-training and evaluation of our on-device LLM.
(a) Outdoor scene (campus walkway)
(b) Indoor scene (corridor)
(c) Data-capture rig (MacBook Pro)
Figure 10: Data-collection environments used for on-device LLM evaluation.